# Linear fractals of the Bezikovitch–Eggleston type

### 03 Mar 2026 v1

@@ 1. Introduction
The latter notation is called the ss–adic representation. We denote by $\alpha_{k}=\alpha_{k}(x)$ the kk–th ss–adic digit of the number $x$. In general, we cannot define the kk–th digit uniquely as a function of the number, since the following equality holds
@@
where $(i)$ denotes the period in the expansion of the number. These numbers are called ss–adic rational and each such number admits exactly two distinct ssadic representations. All other numbers admit a unique expansion, and we call them ss–adic irrational. We agree to use only the first ss–adic representation, namely the one with period $(0)$. Under this agreement, the function $\alpha_{n}(x)$ is well defined on the interval $[0,1]$.
@@
It is clear that the frequency function $\nu_{i}(x)$ of the digit $i$ in the ss–adic representation of a number $x\in[0,1]$ is well defined for ss–adic irrational numbers, and for ss–adic rational numbers, it becomes well defined once we adopt the convention above to use only the representation with period $(0)$.
Since the beginning of the 20th century researchers have actively studied the properties of sets of real numbers by analyzing the frequencies of digits in their representations in various numeral systems. É. Borel and H. Lebesgue produced pioneering results on normal and weakly normal numbers and A. S. Besicovitch, G. Eggleston, P. Billingsley, O. Y. Khinchin, M. M. Postnikov, I. J. Pyatetskii–Shapiro, M. V. Pratsovity, L. Olsen, G. M. Torbin and others investigated properties of sets defined through the frequencies of symbols or combinations of symbols in specific representations.
@@
We call such numbers weakly normal or normal in base $s$. The set of non-normal numbers (those that are not normal) has Lebesgue measure zero and, therefore, may exhibit a fractal structure. A continuous class of fractal subsets of this set arises from the Besicovitch–Eggleston sets, defined as
@@
The asymptotic mean of the digits serves as a certain analogue of the frequency of an ss–adic digit. Moreover, for $s=2$ the equality $r(x)=\nu_{1}(x)$ holds. Clearly, this equality also holds for any number whose ss–adic representation contains only the digits ‘‘0’’ and ‘‘1’’ or even a finite number of other digits. Therefore, for such numbers, the absence of a frequency for the digit ‘‘1’’ (and one can easily construct an example of such a number) is equivalent to the absence of the asymptotic mean of the digits.
@@
Here, the constant $a$ represents a given parameter from $[0,2]$, and we are interested in the topological, metric, and fractal properties of these sets. In particular, in this paper we study the topological, metric, and fractal properties of sets of numbers represented in the ternary system that have a prescribed asymptotic mean of digits and for which the frequencies of all ternary digits exist.
@@ 3. Frequency function of a ternary digit
1. Since $k^{-1}N_{i}(x,k)\leq 1$, the function $\nu_{i}(x)$ cannot take values greater than 1, and therefore it is bounded.
Let us specify a number $x\in[0,1]$ for which $\nu_{i}(x)=a\in[0,1]$. If $a=1$, such a number is $x=\Delta^{3}_{(i)}$, where $i\in{0,1,2}$. If $a=0$, then $x=\Delta^{3}_{(j)}$, where $j\in{0,1,2}$ and $j\neq i$.
Now let $0<a<1$. If $a$ is a rational number, that is $a=\displaystyle\frac{l}{m},$ then the frequency of the ii–th ternary digit of the number
$x=\Delta^{3}_{(\underbrace{i\ldots i}_{l}\underbrace{j\ldots j}_{m-l})},$
is equal to $a$.
@@
where $j\in\{0,1,2\}\setminus\{i\}$. Since $a-\displaystyle\frac{1}{n}=\displaystyle\frac{n\cdot a-1}{n}<\displaystyle\frac{N_{i}(x_{*},n)}{n}=\displaystyle\frac{[n\cdot a]}{n}\leqslant\displaystyle\frac{n\cdot a}{n}=a$ then
@@
Since the frequency does not depend on any finite number of initial digits in a number’s ternary representation, every number in the tail set $x^{*}$ lacks the frequency of both digits $i$ and $j$. Because this tail set is dense in $[0,1]$, the digit frequency function $\nu_{i}(x)$ exhibits a discontinuity within every arbitrarily small interval.
3. If the frequency $\nu_{i}(x_{0})$ does not exist, it is clear that the function $\nu_{i}$ is discontinuous at the point $x_{0}$. Since the definition of continuity at a point requires the function to be defined in some neighborhood of that point, it follows from Proposition 2 that the function $\nu_{i}$ is discontinuous at every point in $[0,1]$. ∎
@@
According to Theorem 1 we have $\nu_{i}(x_{0})=\frac{1}{3}$. It is clear that $x_{n}\to x_{0}$ $(n\to\infty)$ but
@@
$\lim\limits_{n\to\infty}\nu_{i}(x^{\prime}_{n})=1=\lim\limits_{x_{n}\to x_{0}}\nu_{i}(x^{\prime}_{n})\neq\nu_{i}(x_{0}),$
Now let us consider the case when the frequency $\nu_{i}(x)$ in the representation of the number $x$ does not exist. Since $x_{n}\to x_{0}$ $(n\to\infty)$ and $x^{\prime}_{n}\to x_{0}$ $(n\to\infty)$ but
@@ 4. The set of numbers with a prescribed asymptotic mean of digits
Then the sum of the lengths of the first $k$ blocks of ones in the number $\widehat{x}$ equals $v_{k}$, while the sums of the lengths of the first $k$ blocks of zeros and twos are, respectively
@@
$\lim\limits_{k\to\infty}\displaystyle\frac{u_{k}}{N(k)}\geqslant\lim\limits_{k\to\infty}\displaystyle\frac{3^{2k-1}+1}{k}=\lim\limits_{x\to\infty}\displaystyle\frac{3^{2x-1}+1}{x}=\lim\limits_{x\to\infty}\displaystyle\frac{2\cdot 3^{2x-1}\ln 3}{1}=\infty$, i
$\lim\limits_{k\to\infty}\displaystyle\frac{w_{k}}{N(k)}\geqslant\lim\limits_{k\to\infty}\displaystyle\frac{3^{2k}}{k}=\lim\limits_{x\to\infty}\displaystyle\frac{3^{2x}}{x}=\lim\limits_{x\to\infty}\displaystyle\frac{2\cdot 3^{2x}\ln 3}{1}=\infty$, то
@@
Therefore, the Besicovitch–Eggleston set $E[\tau_{0},\tau_{1},\tau_{2}]$ has the cardinality of the continuum.
Since $E[\tau_{0},\tau_{1},\tau_{2}]\subseteq K_{a}$ then the set $K_{a}$ is a continuum set.
@@
From the conditions $\nu_{0}(x)+\nu_{1}(x)+\nu_{2}(x)=1$ and $\nu_{1}(x)+2\nu_{2}(x)=a$ we determine $\nu_{1}=a-2\nu_{2},\>\>\nu_{0}=1-\nu_{1}-\nu_{2}=1-a+2\nu_{2}-\nu_{2}=1-a+\nu_{2}.$ To compute the Hausdorff–Besicovitch dimension, we analyze the function to find its maximum.
@@
$3x^{2}-(1+3a)x+a^{2}=0,\\ x_{1}=\displaystyle\frac{1+3a-\sqrt{1+6a-3a^{2}}}{6}\,\,\text{--- point of maximum},\\ x_{2}=\displaystyle\frac{1+3a+\sqrt{1+6a-3a^{2}}}{6}\,\,\text{---point of minimum},\\$

### 05 Mar 2026 v2

# Linear fractals of the Besicovitch–Eggleston type

@@ 1. Introduction
The latter notation is called the ss–adic representation. We denote by $\alpha_{k}=\alpha_{k}(x)$ the kk–th ss–adic digit of the number $x$. In general we cannot define the kk–th digit uniquely as a function of the number, since the following equality holds
@@
where $(i)$ denotes the period in the expansion of the number. These numbers are called ss–adic rational and each such number admits exactly two distinct ss–adic representations. All other numbers admit a unique expansion, and we call them ss–adic irrational. We agree to use only the first ss–adic representation, namely the one with period $(0)$. Under this agreement, the function $\alpha_{n}(x)$ is well defined on the interval $[0,1]$.
@@
The frequency function $\nu_{i}(x)$ of the digit $i$ in the ss–adic representation of a number $x\in[0,1]$ is well defined for ss–adic irrational numbers, and for ss–adic rational numbers, it becomes well defined once we adopt the convention above to use only the representation with period $(0)$.
Since the beginning of the 20th century researchers have actively studied the properties of sets of real numbers by analyzing the frequencies of digits in their representations in various numeral systems. É. Borel and H. Lebesgue produced pioneering results on normal and weakly normal numbers and A. S. Besicovitch, G. Eggleston, P. Billingsley, O. Y. Khinchin, M. M. Postnikov, I. J. Pyatetskyi–Shapiro, M. V. Pratsiovytyi, L. Olsen, G. M. Torbin and others investigated properties of sets defined through the frequencies of symbols or combinations of symbols in specific representations.
@@
We call such numbers weakly normal or normal on the base $s$. The set of non-normal numbers (those that are not normal) has Lebesgue measure zero and, therefore, may exhibit a fractal structure. A continuous class of fractal subsets of this set arises from the Besicovitch–Eggleston sets, defined as
@@
The asymptotic mean of the digits serves as a certain analogue of the frequency of an ss–adic digit. Moreover, for $s=2$ the equality $r(x)=\nu_{1}(x)$ holds. This equality also holds for any number whose ss–adic representation contains only the digits ‘‘0’’ and ‘‘1’’ or even a finite quantity of other digits. Therefore, for such numbers the absence of a frequency for the digit ‘‘1’’ (and one can easily construct an example of such a number) is equivalent to the absence of the asymptotic mean of the digits.
@@
Here, the constant $a$ represents a given parameter from $[0,2]$, and we are interested in the topological, metric, and fractal properties of these sets. In particular, in this paper we study the topological, metric, and fractal properties of sets of numbers represented in the ternary (33–adic)numeral system that have a prescribed asymptotic mean of digits and for which the frequencies of all ternary digits exist.
@@ 3. Frequency function of a ternary digit
1. Since $k^{-1}N_{i}(x,k)\leq 1$, then the function $\nu_{i}(x)$ cannot take values greater than 1, and therefore it is bounded.
Let us specify a number $x\in[0,1]$ for which $\nu_{i}(x)=a\in[0,1]$. If $a=1$, then such a number is $x=\Delta^{3}_{(i)}$, where $i\in{0,1,2}$. If $a=0$, then $x=\Delta^{3}_{(j)}$, where $j\in{0,1,2}$ and $j\neq i$.
Now let $0<a<1$. If $a$ is a rational number, that is $a=\displaystyle\frac{l}{m},$ then the frequency of the ii–th ternary digit of the number $x=\Delta^{3}_{(\underbrace{i\ldots i}_{l}\underbrace{j\ldots j}_{m-l})},$ is equal to $a$.
@@
and $j\in\{0,1,2\}\setminus\{i\}$. Since $a-\displaystyle\frac{1}{n}=\displaystyle\frac{n\cdot a-1}{n}<\displaystyle\frac{N_{i}(x_{*},n)}{n}=\displaystyle\frac{[n\cdot a]}{n}\leqslant\displaystyle\frac{n\cdot a}{n}=a$, then
@@
Since the frequency does not depend on any finite number of initial digits in a number’s ternary representation, then every number in the tail set $x^{*}$ lacks the frequency of both digits $i$ and $j$. Since this tail set is dense in $[0,1]$, then the digit frequency function $\nu_{i}(x)$ exhibits a discontinuity within every arbitrarily small interval.
3. If the frequency $\nu_{i}(x_{0})$ does not exist, then the function $\nu_{i}$ is discontinuous at the point $x_{0}$. Since the definition of continuity at a point requires the function to be defined in some neighborhood of that point, then it follows from Proposition 2 that the function $\nu_{i}$ is discontinuous at every point in $[0,1]$. ∎
@@
According to Theorem 1 we have $\nu_{i}(x_{0})=\frac{1}{3}$. It is clear that $x_{n}\to x_{0}$ as $n\to\infty$ but
@@
$\lim\limits_{n\to\infty}\nu_{i}(x^{\prime}_{n})=1=\lim\limits_{x_{n}\to x_{0}}\nu_{i}(x^{\prime}_{n})\neq\nu_{i}(x_{0}).$
Now let us consider the case when the frequency $\nu_{i}(x)$ in the representation of the number $x$ does not exist. Since $x_{n}\to x_{0}$ as $n\to\infty$ and $x^{\prime}_{n}\to x_{0}$ as $n\to\infty$ but also
@@ 4. The set of numbers with a prescribed asymptotic mean of digits
Then the sum of the lengths of the first $k$ blocks of ones in the number $\widehat{x}$ equals $v_{k}$, while the sums of the lengths of the first $k$ blocks of zeros and twos are, respectively equals to
@@
$\lim\limits_{k\to\infty}\displaystyle\frac{u_{k}}{N(k)}\geqslant\lim\limits_{k\to\infty}\displaystyle\frac{3^{2k-1}+1}{k}=\lim\limits_{x\to\infty}\displaystyle\frac{3^{2x-1}+1}{x}=\lim\limits_{x\to\infty}\displaystyle\frac{2\cdot 3^{2x-1}\ln 3}{1}=\infty$, and
$\lim\limits_{k\to\infty}\displaystyle\frac{w_{k}}{N(k)}\geqslant\lim\limits_{k\to\infty}\displaystyle\frac{3^{2k}}{k}=\lim\limits_{x\to\infty}\displaystyle\frac{3^{2x}}{x}=\lim\limits_{x\to\infty}\displaystyle\frac{2\cdot 3^{2x}\ln 3}{1}=\infty$, then
@@
Therefore, the Besicovitch–Eggleston set $E[\tau_{0},\tau_{1},\tau_{2}]$ has the cardinality of the continuum. Since $E[\tau_{0},\tau_{1},\tau_{2}]\subseteq K_{a}$ then the set $K_{a}$ is a continuum set.
@@
From the conditions $\nu_{0}(x)+\nu_{1}(x)+\nu_{2}(x)=1$ and $\nu_{1}(x)+2\nu_{2}(x)=a$ we determine $\nu_{1}=a-2\nu_{2}$ and $\nu_{0}=1-\nu_{1}-\nu_{2}=1-a+2\nu_{2}-\nu_{2}=1-a+\nu_{2}.$ To compute the Hausdorff–Besicovitch dimension, we analyze the function to find its maximum.
@@
[…]